I am trying to help a step-grandson and it has been almost 50 years ago I had Algebra II.
Question: Explain how h and k in y=e^(x-h)+ k affects the graph of the exponential function y=e^x

Answer by jim_thompson5910(35256)

Comparing y=e^(x-h)+ k to y=e^x, the 'k' tells us how the vertical translation is done. For example, if k = 2, then this means translate y = e^x two units up. If k = -5, then this tells us to translate 5 units down.

The value of h tells you how to perform the horizontal translation. So if h = 3 for example, then this means translate 3 units to the right. If h = -10, then we translate 10 units to the left.